1/1

Lord George Gordon Byron ,

I Saw Thee Weep

I saw thee weep — the big bright tear Came o’er that eye of blue; And then methought it did appear A violet dropping dew— I saw thee smile — the sapphire’s blaze Beside thee ceased to shine; It could not match the living rays That fill’d that glance of thine.

As clouds from yonder sun receive A deep and mellow dye, Which scarce the shade of coming eve Can banish from the sky— Those smiles unto the moodiest mind Their own pure joy impart; Their sunshine leaves a glow behind That lightens o’er the heart.

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1115-2251 (2015-01-17 13:46:16)

Tiu ĉi poemo troviĝas interalie en http: // poetry. about. com/ od/ poems/ l/ blbyronisawtheeweep. htm .

http://www.poezio.net/version?poem-id=1115&version-id=2251